OpenStudy (richyw):

Show that \[\frac{1}{1/z}=z\]for \(z\neq 0\)

OpenStudy (richyw):

I don't understand how I can show this? \[\left(\left( x+iy \right)^{-1}\right)^{-1}\]

OpenStudy (richyw):

\[\left(\left( x+iy \right)^{-1}\right)^{-1}=x+iy=z\]

OpenStudy (anonymous):

Put \[ z=e^{ i \theta}\\ \frac 1 z = e^{- i \theta}\\ \frac{1}{\frac 1 z}= \frac 1{e^{-i \theta}}=e^{i\theta}=z \]

OpenStudy (richyw):

ok. why do I have to do it in exponential form though?

OpenStudy (anonymous):

The inverse of an inverse of a non zero complex number is the number itself

OpenStudy (anonymous):

You do not need to. That is another way of seeing it

OpenStudy (richyw):

oh ok. just wasn't sure if this was a trick question or something? thanks.